Count the mass of water in which dissolved 250g sucrose will make up 55% sucrose solution?
1
Expert's answer
2020-09-17T07:17:36-0400


mass of solution= mass of sucrose+ mass of water


find the mass of the solution:


"\omega=\frac{ mass of sucrose}{mass of solution}"


"0.55=\frac{250}{mass of solution}"


mass of solution=454.55


mass of water=mass of solution-mass of sucrose=454.55-250=204.55